# Perfect state transfer between real pure states

### v1

@@ 7 Paths
Let $|\sigma_{\mathbf{x}}(A)|\geq 3$ and $\sigma_{\mathbf{x}}(A)$ be closed under algebraic conjugates. As $|\mu_{j}|<2$, Corollary 9 yields $|\sigma_{\mathbf{x}}(A)|=3$. By Lemma 8, we get two cases. If $\sigma_{\mathbf{x}}(A)\subseteq\mathbb{Z}$, then using the same argument as Case 1 of the proof of Theorem 35 to show (1a). If $\sigma_{\mathbf{x}}(A)=\{\frac{a}{2},\frac{1}{2}(a\pm b\sqrt{\Delta})\}$, then one may use the argument in Case 2 to obtain (1b) and (2). ∎
@@
Let $n+1=4m$ and consider $\mathbf{x}=\sum_{j=0}^{\lfloor\frac{m-1}{2}\rfloor}\mathbf{e}_{8j+1}-\sum_{j=1}^{\lfloor\frac{m}{2}\rfloor}\mathbf{e}_{8j-1}$ in $P_{n}$. Observe that we can write $\mathbf{x}=\frac{1}{2}(\mathbf{z}_{2m}+\frac{1}{\sqrt{2}}(\mathbf{z}_{m}+\mathbf{z}_{3m}))$. Invoking Theorem 38(2) with $a=\frac{1}{2}$ and $b=c=\frac{1}{2\sqrt{2}}$, we get PST between $\mathbf{x}$ and $\mathbf{y}=\frac{1}{2}(-\mathbf{z}_{2m}+\frac{1}{\sqrt{2}}(\mathbf{z}_{m}+\mathbf{z}_{3m}))=\sum_{j=0}^{\lfloor\frac{m-1}{2}\rfloor}\mathbf{e}_{8j+3}-\sum_{j=1}^{\lfloor\frac{m}{2}\rfloor}\mathbf{e}_{8j-3}$ at time $\frac{\pi}{\sqrt{2}}$. In particular, observe the following:
@@
Pair perfect state transfer occurs in $P_{n}$ relative to $A$ if and only if $n\in\{3,4,5,7\}$.
@@
Case 1. Let $|\sigma_{\mathbf{e}_{u}-\mathbf{e}_{w}}(A)|=2$. That is, $\mathbf{e}_{u}-\mathbf{e}_{w}=a\mathbf{z}_{j}+b\mathbf{z}_{k}$ and $\mathbf{e}_{v}-\mathbf{e}_{x}=a\mathbf{z}_{j}-b\mathbf{z}_{k}$ with $j\neq k$. This implies that $\sin(\frac{j\ell\pi}{n+1})=\sin(\frac{k\ell\pi}{n+1})=0$ for any $\ell\neq u,w,v,x$. Since $j\neq k$, we get $\ell\neq 1,2$, so we may assume that $u=1$ and either $v=2$ or $w=2$. By Lemma 5, if $|\sigma_{\mathbf{e}_{1}-\mathbf{e}_{w}}(A)|=2$, then the covering radius of $\mathbf{e}_{1}-\mathbf{e}_{w}$ is one. Hence, $n=3$ and $w=2$ or $n\in\{4,5\}$ and $w=n$. Using Lemma 21, one checks that PST occurs between (i) $\mathbf{e}_{1}-\mathbf{e}_{2}$ and $\mathbf{e}_{3}-\mathbf{e}_{2}$ in $P_{3}$ at time $\frac{\pi}{\sqrt{2}}$, (ii) $\mathbf{e}_{1}-\mathbf{e}_{4}$ and $\mathbf{e}_{2}-\mathbf{e}_{3}$ in $P_{4}$ at time $\frac{\pi}{\sqrt{5}}$, (iii) $\mathbf{e}_{1}-\mathbf{e}_{5}$ and $\mathbf{e}_{2}-\mathbf{e}_{4}$ in $P_{5}$ at time $\pi$.
@@
An analogous argument yields all cases of plus PST in $P_{n}$: between $\mathbf{e}_{1}+\mathbf{e}_{2}$ and $\mathbf{e}_{3}+\mathbf{e}_{2}$ in $P_{3}$ at time $\frac{\pi}{\sqrt{2}}$, $\mathbf{e}_{1}+\mathbf{e}_{4}$ and $\mathbf{e}_{2}+\mathbf{e}_{3}$ in $P_{4}$ at time $\frac{\pi}{\sqrt{5}}$, $\mathbf{e}_{1}+\mathbf{e}_{5}$ and $\mathbf{e}_{2}+\mathbf{e}_{4}$ in $P_{5}$ at time $\frac{\pi}{\sqrt{3}}$.
@@
Plus perfect state transfer occurs in $P_{n}$ relative to $A$ if and only if $n\in\{3,4,5\}$.
@@
We now turn to to the Laplacian case. The Laplacian eigenvalues and eigenvectors of $P_{n}$ are known, see [brouwer2011spectra, Section 1.4.3]. We provide normalized eigenvectors for $L(P_{n})$ below.
@@
    $n=3m$, $\mathbf{x}=a\mathbf{w}_{\frac{3m}{2}}+b\mathbf{w}_{\frac{m}{2}}+c\mathbf{w}_{\frac{5m}{2}}$ and $\mathbf{y}=a\mathbf{w}_{\frac{3m}{2}}-b\mathbf{w}_{\frac{m}{2}}-c\mathbf{w}_{\frac{5m}{2}}$.
@@
    In $P_{3}$, $\{\mathbf{e}_{1}-\mathbf{e}_{2},\mathbf{e}_{3}-\mathbf{e}_{2}\}$ admits PST at $\frac{\pi}{2}$. Moreover, $\mathbf{e}_{1}+\mathbf{e}_{2}$ and $\mathbf{e}_{3}+\mathbf{e}_{2}$ are strongly cospectral and periodic with $\sigma_{\mathbf{x},\mathbf{y}}^{+}(L)=\{0,3\}$ and $\sigma_{\mathbf{x},\mathbf{y}}^{-}(L)=\{1\}$, but they do not admit PST as Corollary 28(3) does not hold.
@@
    In $P_{4}$, $\{\mathbf{e}_{1}+\mathbf{e}_{4},\mathbf{e}_{2}+\mathbf{e}_{3}\}$ admits PST at $\frac{\pi}{2}$, while the pairs $\{\mathbf{e}_{1}-\mathbf{e}_{2},\mathbf{e}_{3}-\mathbf{e}_{4}\}$, $\{\mathbf{e}_{1}+\mathbf{e}_{4},\mathbf{e}_{2}+\mathbf{e}_{3}\}$, $\{\mathbf{e}_{2}-\mathbf{e}_{3},\frac{1}{\sqrt{2}}(\mathbf{e}_{1}-\mathbf{e}_{2}+\mathbf{e}_{3}-\mathbf{e}_{4})\}$ and $\{\mathbf{e}_{1}-\mathbf{e}_{4},\frac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2}-\mathbf{e}_{3}-\mathbf{e}_{4})\}$ all admit PST at $\frac{\pi}{\sqrt{2}}$.
@@
    In $P_{5}$, $\{\mathbf{e}_{1}-\mathbf{e}_{5},\frac{1}{\sqrt{5}}(\mathbf{e}_{1}+2\mathbf{e}_{2}-2\mathbf{e}_{4}-\mathbf{e}_{5})\}$ and $\{\mathbf{e}_{2}-\mathbf{e}_{4},\frac{1}{\sqrt{5}}(2\mathbf{e}_{1}-\mathbf{e}_{2}+\mathbf{e}_{4}-2\mathbf{e}_{5})\}$ admit PST, both at $\frac{\pi}{\sqrt{5}}$.
@@ 8 Cartesian product
Let $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{m}$ and $\mathbf{y}_{1},\mathbf{y}_{2}\in\mathbb{R}^{n}$ such that $\mathbf{x}_{1}\neq\mathbf{y}_{1}$. Then perfect state transfer occurs between $\mathbf{x}_{1}\otimes\mathbf{y}_{1}$ and $\mathbf{x}_{2}\otimes\mathbf{y}_{2}$ in $G\square H$ at time $\tau$ if and only if either
@@
  • By Corollary 48(1) and Example 36, $Q_{d}\square C_{8}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}+\mathbf{e}_{5}),\mathbf{e}_{v}\otimes(\mathbf{e}_{3}+\mathbf{e}_{7})\}$, and $Q_{d}\square C_{12}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}+\mathbf{e}_{5}+\mathbf{e}_{9}),\mathbf{e}_{v}\otimes(\mathbf{e}_{3}+\mathbf{e}_{7}+\mathbf{e}_{11})\}$ at $\frac{\pi}{2}$.
  • By Corollary 48 and [kim2024generalization, Theorem 6.5(iv-vi)], $Q_{d}\square C_{6}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}-\mathbf{e}_{3}),\mathbf{e}_{v}\otimes(\mathbf{e}_{4}-\mathbf{e}_{6})\}$ at $\frac{\pi}{2}$, and between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}+2\mathbf{e}_{3}),\mathbf{e}_{u}\otimes(\mathbf{e}_{1}+2\mathbf{e}_{5})\}$ and $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}+\frac{1}{2}\mathbf{e}_{3}),\mathbf{e}_{u}\otimes(\mathbf{e}_{5}+\frac{1}{2}\mathbf{e}_{3})\}$ at $\pi$.
@@
Let $M=A$ and consider $P_{3}^{\square n}$, the Cartesian product of $n\geq 1$ copies of $P_{3}$. This graph admits PST at time $\frac{\pi}{\sqrt{2}}$ between any pair of vertices $\mathbf{e}_{u}$ and $\mathbf{e}_{v}$ at distance $2n$. By Corollary 48(1) and Example 39(1), $(P_{3}^{\square n})\square P_{7}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}-\mathbf{e}_{7}),\mathbf{e}_{v}\otimes(\mathbf{e}_{3}-\mathbf{e}_{5})\}$ at $\frac{\pi}{\sqrt{2}}$ for all $n\geq 1$. Moreover, by Corollary 48(1) with [kim2024generalization, Theorem 6.5(vii)], $(P_{3}^{\square n})\square C_{8}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}-\mathbf{e}_{3}),\mathbf{e}_{v}\otimes(\mathbf{e}_{5}-\mathbf{e}_{7})\}$ at $\frac{\pi}{2}$ for all $n\geq 1$.
@@ 9 Joins
Let $G$ and $H$ be weighted graphs on $m$ and $n$ vertices, respectively. The join of $G$ and $H$ is obtained from taking a copy of $G$ and a copy of $H$ and adding all edges between $G$ and $H$ with weight one. Throughout, we assume that $G$ and $H$ are $k$- and $\ell$-regular graphs, respectively whenever we deal with $M=A$.
@@
whenever $G$ and $H$ are connected. If $G$ (respectively, $H$) is disconnected, then we include the term $e^{itn}\left[\begin{array}[]{cc}E_{0}-\frac{1}{m}J&{\bf{0}}\\ {\bf{0}}&{\bf{0}}\end{array}\right]$ (respectively, $e^{itm}\left[\begin{array}[]{cc}{\bf{0}}&{\bf{0}}\\ {\bf{0}}&F_{0}-\frac{1}{n}J\end{array}\right]$) in the third (respectively, fourth) summand in (8).
@@
    Suppose $G$ admits perfect state transfer between $\mathbf{x}_{1}$ and $\mathbf{y}_{1}$ and $H$ admits perfect state transfer between $\mathbf{x}_{2}$ and $\mathbf{y}_{2}$ both at time $\tau$. If $\tau(\lambda-\theta+\delta(n-m))\equiv 0$ (mod $2\pi$) for some $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(A)$, $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(A)$ and $\delta\in\{0,1\}$, then $G\vee H$ admits perfect state transfer between $\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]$ and $\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ at time $\tau$ relative to $M=L$ whenever $\delta=1$ and relative to $M=A$ whenever $\delta=0$.
@@
Let $U_{G}(t)$ denote the transition matrix of $G$ relative to $A$ or $L$. As ${\bf{1}}^{\top}\mathbf{x}_{1}=0$, equation (8) yields
@@
As PST occurs between $\mathbf{x}_{j}$ and $\mathbf{y}_{j}$ for $j\in\{1,2\}$ at $\tau$, we have $U_{G}(\tau)\mathbf{x}_{1}=\gamma_{1}\mathbf{y}_{1}$ and $U_{H}(\tau)\mathbf{x}_{2}=\gamma_{2}\mathbf{y}_{2}$, where $\gamma_{1}=e^{i\tau\lambda}$ and $\gamma_{2}=e^{i\tau\theta}$ for all $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(M)$ and $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(M)$. Thus, if $\tau(\lambda-\theta+n-m)\equiv 0$ (mod $2\pi$) for some $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(M)$ and $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(M)$, then $\tau(n+\lambda)\equiv\tau(m+\theta)$ (mod $2\pi$), and so $e^{i\tau(n+\lambda)}=e^{i\tau(m+\theta)}$. Thus, $U_{L}(\tau)\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]=\left[\begin{array}[]{cc}e^{i\tau n}U_{G}(\tau)\mathbf{x}_{1}\\ e^{i\tau m}U_{H}(\tau)\mathbf{x}_{2}\end{array}\right]=\left[\begin{array}[]{cc}e^{i\tau(n+\lambda)}\gamma_{1}\mathbf{y}_{1}\\ e^{i\tau(m+\theta)}\mathbf{y}_{2}\end{array}\right]=e^{i\tau(n+\lambda)}\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ by equation (10). This proves that PST occurs between $\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]$ and $\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ in $G\vee H$. The same argument applies to $M=A$ except that $e^{itn}$ and $e^{itm}$ in (10) are both absent. ∎

### v2

@@ 7 Paths
Let $|\sigma_{\mathbf{x}}(A)|\geq 3$ and $\sigma_{\mathbf{x}}(A)$ be closed under algebraic conjugates. As $|\mu_{j}|<2$, Corollary 9 yields $|\sigma_{\mathbf{x}}(A)|=3$. By Theorem 8, we get two cases. If $\sigma_{\mathbf{x}}(A)\subseteq\mathbb{Z}$, then using the same argument as Case 1 of the proof of Theorem 35 to show (1a). If $\sigma_{\mathbf{x}}(A)=\{\frac{a}{2},\frac{1}{2}(a\pm b\sqrt{\Delta})\}$, then one may use the argument in Case 2 to obtain (1b) and (2). ∎
@@
Let $n+1=4m$ and consider $\mathbf{x}=\sum_{j=0}^{\lfloor\frac{m-1}{2}\rfloor}\mathbf{e}_{8j+1}-\sum_{j=1}^{\lfloor\frac{m}{2}\rfloor}\mathbf{e}_{8j-1}$ in $P_{n}$. Observe that we can write $\mathbf{x}=\frac{\sqrt{n+1}}{2\sqrt{2}}(\mathbf{z}_{2m}+\frac{1}{\sqrt{2}}(\mathbf{z}_{m}+\mathbf{z}_{3m}))$. Invoking Theorem 38(2) with $a=\sqrt{2}b=\sqrt{2}c=\frac{\sqrt{n+1}}{2\sqrt{2}}$, we get PST between $\mathbf{x}$ and $\mathbf{y}=\frac{\sqrt{n+1}}{2\sqrt{2}}(-\mathbf{z}_{2m}+\frac{1}{\sqrt{2}}(\mathbf{z}_{m}+\mathbf{z}_{3m}))=\sum_{j=0}^{\lfloor\frac{m-1}{2}\rfloor}\mathbf{e}_{8j+3}-\sum_{j=1}^{\lfloor\frac{m}{2}\rfloor}\mathbf{e}_{8j-3}$ at time $\frac{\pi}{\sqrt{2}}$. In particular, the following hold.
@@
Pair perfect state transfer occurs in $P_{n}$ relative to $A$ if and only if $n\in\{3,5,7\}$.
@@
Case 1. Let $|\sigma_{\mathbf{e}_{u}-\mathbf{e}_{w}}(A)|=2$. That is, $\mathbf{e}_{u}-\mathbf{e}_{w}=a\mathbf{z}_{j}+b\mathbf{z}_{k}$ and $\mathbf{e}_{v}-\mathbf{e}_{x}=a\mathbf{z}_{j}-b\mathbf{z}_{k}$ with $j\neq k$. This implies that $\sin(\frac{j\ell\pi}{n+1})=\sin(\frac{k\ell\pi}{n+1})=0$ for any $\ell\neq u,w,v,x$. Since $j\neq k$, we get $\ell\neq 1,2$, so we may assume that $u=1$ and either $v=2$ or $w=2$. By Lemma 5, if $|\sigma_{\mathbf{e}_{1}-\mathbf{e}_{w}}(A)|=2$, then the covering radius of $\mathbf{e}_{1}-\mathbf{e}_{w}$ is one. Hence, $n=3$ and $w=2$ or $n\in\{4,5\}$ and $w=n$. Using Lemma 21, one checks that $\mathbf{e}_{1}-\mathbf{e}_{4}$ and $\mathbf{e}_{2}-\mathbf{e}_{3}$ are not strongly cospectral in $P_{4}$, while PST occurs between $\mathbf{e}_{1}-\mathbf{e}_{2}$ and $\mathbf{e}_{3}-\mathbf{e}_{2}$ in $P_{3}$ at $\frac{\pi}{\sqrt{2}}$. In [pal2024quantum], it was observed that $P_{5}$ has PST between $\mathbf{e}_{1}-\mathbf{e}_{5}$ and $\mathbf{e}_{2}-\mathbf{e}_{4}$ at $\frac{\pi}{2}$.
@@
An analogous argument yields $P_{3}$ as the only path that admits PST, between $\mathbf{e}_{1}+\mathbf{e}_{2}$ and $\mathbf{e}_{3}+\mathbf{e}_{2}$ in $P_{3}$ at $\frac{\pi}{\sqrt{2}}$.
@@
Plus perfect state transfer occurs in $P_{n}$ relative to $A$ if and only if $n=3$.
@@
We now turn to the Laplacian case. The Laplacian eigenvalues and eigenvectors of $P_{n}$ are known, see [brouwer2011spectra, Section 1.4.3]. We provide normalized eigenvectors for $L(P_{n})$ below.
@@
    $n=6m$, $\mathbf{x}=a\mathbf{w}_{\frac{3m}{2}}+b\mathbf{w}_{\frac{m}{2}}+c\mathbf{w}_{\frac{5m}{2}}$ and $\mathbf{y}=a\mathbf{w}_{\frac{3m}{2}}-b\mathbf{w}_{\frac{m}{2}}-c\mathbf{w}_{\frac{5m}{2}}$.
@@
    In $P_{3}$, $\{\mathbf{e}_{1}-\mathbf{e}_{2},\mathbf{e}_{3}-\mathbf{e}_{2}\}$ has PST at $\frac{\pi}{2}$. Moreover, $\mathbf{e}_{1}+\mathbf{e}_{2}$ and $\mathbf{e}_{3}+\mathbf{e}_{2}$ are strongly cospectral and periodic with $\sigma_{\mathbf{x},\mathbf{y}}^{+}(L)=\{0,3\}$ and $\sigma_{\mathbf{x},\mathbf{y}}^{-}(L)=\{1\}$, but they do not admit PST as Corollary 28(3) does not hold.
@@
    In $P_{4}$, $\{\mathbf{e}_{1}+\mathbf{e}_{4},\mathbf{e}_{2}+\mathbf{e}_{3}\}$ has PST at $\frac{\pi}{2}$, $\{\mathbf{e}_{1}-\mathbf{e}_{2},\mathbf{e}_{3}-\mathbf{e}_{4}\}$ has PST at $\frac{\pi}{\sqrt{2}}$, and $\{\mathbf{e}_{2}-\mathbf{e}_{3},\frac{1}{\sqrt{2}}(\mathbf{e}_{1}-\mathbf{e}_{2}+\mathbf{e}_{3}-\mathbf{e}_{4})\}$ and $\{\mathbf{e}_{1}-\mathbf{e}_{4},\frac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2}-\mathbf{e}_{3}-\mathbf{e}_{4})\}$ have PST at $\frac{\pi}{2\sqrt{2}}$.
@@
    In $P_{5}$, $\{\mathbf{e}_{1}-\mathbf{e}_{5},\frac{1}{\sqrt{5}}(\mathbf{e}_{1}+2\mathbf{e}_{2}-2\mathbf{e}_{4}-\mathbf{e}_{5})\}$ and $\{\mathbf{e}_{2}-\mathbf{e}_{4},\frac{1}{\sqrt{5}}(2\mathbf{e}_{1}-\mathbf{e}_{2}+\mathbf{e}_{4}-2\mathbf{e}_{5})\}$ have PST, both at $\frac{\pi}{\sqrt{5}}$.
@@ 8 Cartesian product
Let $\mathbf{x}_{1},\mathbf{y}_{1}\in\mathbb{R}^{m}$ and $\mathbf{x}_{2},\mathbf{y}_{2}\in\mathbb{R}^{n}$ such that $\mathbf{x}_{1}\neq\mathbf{y}_{1}$. Then perfect state transfer occurs between $\mathbf{x}_{1}\otimes\mathbf{x}_{2}$ and $\mathbf{y}_{1}\otimes\mathbf{y}_{2}$ in $G\square H$ at time $\tau$ if and only if either
@@
  • By Corollary 48(1) and Example 36, $Q_{d}\square C_{8}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}+\mathbf{e}_{4}),\mathbf{e}_{v}\otimes(\mathbf{e}_{2}+\mathbf{e}_{6})\}$, and $Q_{d}\square C_{12}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}+\mathbf{e}_{4}+\mathbf{e}_{8}),\mathbf{e}_{v}\otimes(\mathbf{e}_{2}+\mathbf{e}_{6}+\mathbf{e}_{10})\}$ at $\frac{\pi}{2}$.
  • By Corollary 48 and [kim2024generalization, Theorem 6.5(iv-vi)], $Q_{d}\square C_{6}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}-\mathbf{e}_{2}),\mathbf{e}_{v}\otimes(\mathbf{e}_{3}-\mathbf{e}_{5})\}$ at $\frac{\pi}{2}$, and between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}+2\mathbf{e}_{2}),\mathbf{e}_{u}\otimes(\mathbf{e}_{0}+2\mathbf{e}_{4})\}$ and $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}+\frac{1}{2}\mathbf{e}_{2}),\mathbf{e}_{u}\otimes(\mathbf{e}_{4}+\frac{1}{2}\mathbf{e}_{2})\}$ at $\pi$.
@@
Let $M=A$ and consider $P_{3}^{\square n}$, the Cartesian product of $n\geq 1$ copies of $P_{3}$. This graph admits PST at time $\frac{\pi}{\sqrt{2}}$ between any pair of vertices $\mathbf{e}_{u}$ and $\mathbf{e}_{v}$ at distance $2n$. By Corollary 48(1) and Example 39(1), $(P_{3}^{\square n})\square P_{7}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{1}-\mathbf{e}_{7}),\mathbf{e}_{v}\otimes(\mathbf{e}_{3}-\mathbf{e}_{5})\}$ at $\frac{\pi}{\sqrt{2}}$ for all $n\geq 1$. Moreover, by Corollary 48(1) with [kim2024generalization, Theorem 6.5(vii)], $(P_{3}^{\square n})\square C_{8}$ admits PST between $\{\mathbf{e}_{u}\otimes(\mathbf{e}_{0}-\mathbf{e}_{2}),\mathbf{e}_{v}\otimes(\mathbf{e}_{4}-\mathbf{e}_{6})\}$ at $\frac{\pi}{\sqrt{2}}$ for all $n\geq 1$.
@@ 9 Joins
Let $G$ and $H$ be weighted graphs on $m$ and $n$ vertices, respectively. The join of $G$ and $H$, denoted $G\vee H$, is obtained from taking a copy of $G$ and a copy of $H$ and adding all edges between $G$ and $H$ with weight one. Throughout, we assume that $G$ and $H$ are $k$- and $\ell$-regular graphs, respectively whenever we deal with $M=A$.
@@
whenever $G$ and $H$ are connected. If $G$ (respectively, $H$) is disconnected, then we add the term $e^{itn}\left[\begin{array}[]{cc}E_{0}-\frac{1}{m}J&{\bf{0}}\\ {\bf{0}}&{\bf{0}}\end{array}\right]$ (respectively, $e^{itm}\left[\begin{array}[]{cc}{\bf{0}}&{\bf{0}}\\ {\bf{0}}&F_{0}-\frac{1}{n}J\end{array}\right]$) in the third (respectively, fourth) summand in (9).
@@
    Suppose $G$ admits perfect state transfer between $\mathbf{x}_{1}$ and $\mathbf{y}_{1}$ and $H$ admits perfect state transfer between $\mathbf{x}_{2}$ and $\mathbf{y}_{2}$ both at time $\tau$. If $\tau(\lambda-\theta+\delta(n-m))\equiv 0$ (mod $2\pi$) for some $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(M)$, $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(M)$ and $\delta\in\{0,1\}$, then $G\vee H$ admits perfect state transfer between $\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]$ and $\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ at time $\tau$ relative to $M=L$ whenever $\delta=1$ and relative to $M=A$ whenever $\delta=0$.
@@
Let $U_{G}(t)$ denote the transition matrix of $G$ relative to $L$. As ${\bf{1}}^{\top}\mathbf{x}_{1}=0$, equation (9) yields
@@
As PST occurs between $\mathbf{x}_{j}$ and $\mathbf{y}_{j}$ for $j\in\{1,2\}$ at $\tau$, we have $U_{G}(\tau)\mathbf{x}_{1}=\gamma_{1}\mathbf{y}_{1}$ and $U_{H}(\tau)\mathbf{x}_{2}=\gamma_{2}\mathbf{y}_{2}$, where $\gamma_{1}=e^{i\tau\lambda}$ and $\gamma_{2}=e^{i\tau\theta}$ for all $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(L)$ and $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(L)$. Thus, if $\tau(\lambda-\theta+n-m)\equiv 0$ (mod $2\pi$) for some $\lambda\in\sigma_{\mathbf{x}_{1},\mathbf{y}_{1}}^{+}(M)$ and $\theta\in\sigma_{\mathbf{x}_{2},\mathbf{y}_{2}}^{+}(M)$, then $\tau(n+\lambda)\equiv\tau(m+\theta)$ (mod $2\pi$), and so $e^{i\tau(n+\lambda)}=e^{i\tau(m+\theta)}$. Thus, $U_{L}(\tau)\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]=\left[\begin{array}[]{cc}e^{i\tau n}U_{G}(\tau)\mathbf{x}_{1}\\ e^{i\tau m}U_{H}(\tau)\mathbf{x}_{2}\end{array}\right]=\left[\begin{array}[]{cc}e^{i\tau(n+\lambda)}\gamma_{1}\mathbf{y}_{1}\\ e^{i\tau(m+\theta)}\mathbf{y}_{2}\end{array}\right]=e^{i\tau(n+\lambda)}\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ by equation (11). This proves that PST occurs between $\left[\begin{array}[]{cc}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{array}\right]$ and $\left[\begin{array}[]{cc}\mathbf{y}_{1}\\ \mathbf{y}_{2}\end{array}\right]$ in $G\vee H$. The same argument applies to $M=A$ except that $e^{itn}$ and $e^{itm}$ in (11) are both absent. ∎